ANDREW P. DEASON

Berwyn, IL / adeason at dson dot org

OVERVIEW

Current freelance software developer and consultant for OpenAFS and related technology. Special interest in computing technology, systems administration, systems programming and related areas.

SKILLS

∘ Software Development: C, git, Perl, Python, C++, SQL, Javascript, and some CVS/svn/Hg ∘ AFS: Cell maintenance, cell architecture design, and OpenAFS feature development and bug fixes ∘ Database Administration: PostgreSQL, MySQL, and SQLite ∘ Network Administration: KerberosV (MIT, Heimdal, some AD), LDAP, and DNS (bind9) ∘ Server Administration: Virtualization (KVM, Xen, Solaris containers, VMWare), and

Windows/Unix integration ∘ Web Administration: Nginx, Apache HTTPd, XSLT, and HTML/XHTML/CSS ∘ Operating Systems: Debian, CentOS/RHEL, Solaris 10 and 11, and some Solaris 8 and 9, AIX 5, and

HP-UX 11i

EXPERIENCE

Contractor – July 2015 - Present Senior Software Engineer – November 2011 - June 2015 Software Engineer – June 2009 - October 2011 OpenAFS Consultant – May 2008 - May 2009 Sine Nomine Associates—http://www.sinenomine.net/

Involved in various OpenAFS-related responsibilities, including but not limited to customer trouble ticket resolution, OpenAFS feature development, bug fixes, documentation, and testing. Worked directly with customers, users, community members, and other developers and maintainers.

Systems Administrator – Fall 2007 - Spring 2009 Calculus & Mathematica at UIUC—http://cm.math.uiuc.edu/

Maintained multiple systems for the Calculus & Mathematica program, including the Classcomm electronic handin application, email, and various frontend websites for the program.

Systems Administrator – January 2007 - Spring 2009 Association for Computing Machinery at UIUC—http://www.acm.uiuc.edu/

Maintained several systems supporting numerous ACM SIG projects and ACM workstations along with a team of systems administrators. Transitioned the environment to LDAP from NIS, maintained the acm.uiuc.edu AFS cell, set up and maintained the INN2 news server, and maintained various Solaris and Linux machines.

Linux Administrator – January 2007 - Spring 2009 UIUC Engineering Council—http://www.ec.uiuc.edu/

Helped transition numerous services from an old Red Hat Linux server to a new Mac OSX machine. Integrated with Windows-hosted file storage and was responsible for setting up and maintaining new Linux systems.

Systems Administrator I – May 2007 - May 2008 Science Applications International Corporation (SAIC)—https://www.saic.com/

Set up and maintained the test environment and infrastructure for the ISMS project, an Oracle-backed J2EE survey application. Created and maintained several Xen paravirtualized environments with custom security enhancements, and automated the interaction between environments.

Systems Administrator – Summer 2004, 2005, and 2006 American Council On Renewable Energy (ACORE)—https://www.acore.org/

Set up and maintained the ACORE Salesforce.com CRM account, and semi-automated the import of thousands of contacts from Microsoft Excel spreadsheets. Helped recreate the ACORE website and transition resources to a new domain name. Set up a modified version of WordPress for the ACORE Renewable Energy Blog.

Head Systems Administrator – June 2005 - June 2006 Webmaster – June 2004 - June 2006 Systems Administrator – December 2003 - June 2006 Thomas Jefferson High School for Science and Technology—https://www.tjhsst.edu/

Managed the TJHSST Computer Systems Labwith a team of fellow student Systems Administrators. Maintained the www.tjhsst.edu website, co-developed the school intranet in PHP5, assited with integrating the Linux-hosted mail system with Active Directory, maintained and helped design the csl.tjhsst.edu AFS cell, and integrated an Exabyte EXB-480 tape drive carousel with AMANDA and OpenAFS’s native backup system.

Junior Developer/Intern – Summer 2004 Interactive Multimedia Applications Group (IMAG)—https://it.gwu.edu/

Created and assisted with the creation of various web sites requested of IMAG (including working on the new portal for the George Washington University, myGW), utilizing ColdFusion MX and Oracle SQL.

PRESENTATIONS

European AFS & Kerberos Conference 2014—https://indico.cern.ch/event/271400

OpenAFS’ Road to IPv6 (co-presenter) – 27 Mar 2014

OOB Update: Pipelines – 26 Mar 2014

European AFS & Kerberos Conference 2012—http://conferences.inf.ed.ac.uk/eakc2012/

OpenAFS Out-of-Band TCP – 16 Oct 2012

European AFS & Kerberos Conference 2011—https://indico.desy.de/event/eakc2011

afsd.fuse – 5 October 2011

AFS & Kerberos Best Practices Workshop—http://workshop.openafs.org/

Real-world AFS Testing with ’afsload’ – 17 June 2011

Deploying the Demand Attach File Server (DAFS) (co-presenter) – 16 June 2011

libuafs: The coolest OpenAFS library you’ve never heard of – 27 May 2010

Jefferson Overseas Schools Technology Institute (JOSTI)—http://information.tjhsst.edu/overseas/ Panel: Open Source Advantages and Limitations (panelist) – 1 July 2004

Samba: Bridging the Gap (assistant) – 1 July 2004

Intro to Unix: A Survival Kit (co-presenter) – 29 June 2004

EDUCATION

University of Illinois at Urbana-Champaign – Fall 2006 - Spring 2009 Urbana, Illinois—http://illinois.edu/

Received a Bachelor of Science in Computer Science. Graduated with High Honors, and a 3.91 GPA.

Thomas Jefferson High School for Science and Technology – September 2002 - June 2006 Alexandria, VA—http://www.tjhsst.edu/

Graduated with a 4.015 GPA. Final project involved recreating the school Intranet application in Object-Oriented PHP5.

Last updated: 24 August 2017 06:24 UTC The most recent version of this resume can be found at: http://www.dson.org/^{∼}adeason/resume/